# UNIMO-G: Unified Image Generation through Multimodal Conditional Diffusion

### v1

@@ Abstract
**footnotetext: These authors contributed equally to this work.
@@ 1 Introduction
Recent advancements in text-to-image (T2I) diffusion models have yielded impressive results in the generation of high-fidelity images from textual descriptions. Various methods, including DALLE-2 Ramesh et al. (2022), Imagen Saharia et al. (2022), Stable Diffusion Rombach et al. (2022), and SD-XL Podell et al. (2023), have been successful in producing photo-realistic and contextually relevant images based on textual prompts. Nevertheless, a fundamental challenge persists due to the inherent brevity of textual descriptions, particularly when intricate details, specific entities, or nuanced scenes are involved. Thus, faithfully generating images from general vision-language (VL) inputs is essential to improve the controllability of image generation.
@@
UNIMO-G exhibits a comprehensive capability for controllable image generation, excelling not only in text-to-image synthesis but also in zero-shot subject-driven generation. It adeptly produces high-fidelity images from multimodal prompts, even those containing multiple image entities. To assess its performance, we conducted evaluations in both text-to-image and subject-driven generation contexts using the MS-COCO and DreamBench datasets, respectively. The results consistently highlight UNIMO-G’s superior performance in these scenarios. Additionally, recognizing DreamBench’s focus on single-subject generation, we introduce MultiBench, a new benchmark featuring images with multiple entities. The evaluation on MultiBench confirms UNIMO-G’s effectiveness in zero-shot multi-entity subject-driven generation.
@@
Figure 2: UNIMO-G consists of an MLLM for multimodal perception, and a conditional denoising UNet for image generation. It accepts multimodal prompts with interleaved images and texts, and generate images consistent with the image entities. Orange denotes the trainable modules; Blue denotes the frozen ones.
@@ 2 Method
The architecture of UNIMO-G, as depicted in Figure 2, primarily comprises two key components: a Multimodal Large Language Model (MLLM) responsible for encoding multimodal prompts and a conditional denoising diffusion network for image generation based on the encoded representations. In our study, we employed an in-house Chinese MLLM, structurally analogous to MiniGPT-4 Zhu et al. (2023). It contains a total of 7 billion parameters, underwent pre-training on a vast dataset comprising billions of image-text pairs. This extensive training process equips the model with a robust capability to process and interpret complex multimodal data. The training of UNIMO-G is conducted in a two-stage process:
@@ 2.1 Text-to-Image Pre-training
$Attn(z_{t},c)=softmax(\frac{Q(z_{t})\cdot K(c)^{T})}{\sqrt{d}})\cdot V(c)$
@@
Training a text-to-image diffusion model from scratch presents significant challenges in terms of complexity and resource expenditure. To address these, we introduce an effective training schedule to enhance the efficiency and performance of model training. This schedule encompasses three phases: (1) initial training on a small image corpus to establish foundational visual distribution knowledge; (2) subsequent expansion to large-scale text-image pair training, focusing on text-visual correspondence; (3) culminating in a final phase of training with a small refined corpus, characterized by high visual aesthetics and precise text-image alignment. In our experiments, the training of the UNet model is initially conducted using the CC3M dataset Sharma et al. (2018). This dataset is chosen for its diverse range of visual concepts coupled with straightforward image descriptions, making it an effective tool for initiating training from scratch. Subsequently, the model undergoes further training with an extensive collection of 300M text-image pairs, aimed at broadening its conceptual understanding and improving its alignment to textual descriptions. The final phase of training involves fine-tuning the model using a meticulously curated corpus, consisting of tens of thousands of high-quality image-text pairs, carefully selected for their superior quality. Based on the above strategies and our architecture designs, we obtain a powerful Chinese text-to-Image generation model, surpassing open-source models like Stable Diffusion and its advanced version SD-XL Podell et al. (2023). We provide detailed evaluations of our results in Section 4.2 and implementation details in Appendix A.
@@ 2.2 Multimodal Instruction Tuning
To enhance the representativeness of text prompts, we introduce a format for multimodal prompts that are composed of interleaved images and texts. Specifically, entities mentioned in text captions can be substituted with their corresponding images, like “<img>image of Elon Musk</img> holding his <img>image of iPhone</img>, takes a selfie on the street”, as shown in Figure 2. To create pairs of multimodal prompts and images, we have designed a data processing pipeline as illustrated in Figure 3. The pipeline first generates captions and extracts entities from the caption by MLLM. Subsequently, it acquires the corresponding image segment for each entity using a combination of language grounding by Grounding DINO Liu et al. (2023) and image segmentation by SAM Kirillov et al. (2023). Further introduction on the data construction process is provided in Section A. With a collection of pairs of multimodal prompts and images, UNIMO-G is trained to generate images in accordance with these multimodal instructions.
@@
$CA(z_{t},v)=Softmax(Q(z_{t})\cdot K(v)^{T})$
@@ 3 Related Work
The incorporation of diffusion models into text-to-image synthesis represents a notable advancement in computational creativity Ho et al. (2020); Song et al. (2020); Li et al. (2022). Models like GLIDE Nichol et al. (2021) and DALL-E 2 Ramesh et al. (2022), which utilize CLIP image embeddings, have substantially progressed in producing images that are both diverse and semantically coherent with textual inputs. Imagen Saharia et al. (2022) underscores the importance of language comprehension, proposing the integration of a large T5 language model to enhance semantic representation. The Latent Diffusion model Rombach et al. (2022) addresses computational constraints by generating images from text-conditioned, low-dimensional latent spaces. Our proposed framework builds upon the principles of the Latent Diffusion model, leveraging its computational efficiency and scalability.
@@
Following the success of generating high quality images from text descriptions, recent studies have explored subject-driven generation techniques. Models like DreamBooth Ruiz et al. (2023), textual-inversion Gal et al. (2022), and custom-diffusion Kumari et al. (2023) use optimization-based methods to embed subjects into diffusion models. This is achieved by either fine-tuning the model weights Ruiz et al. (2023); Kumari et al. (2023) or inverting the subject image into a text token that encodes the subject identity Gal et al. (2022). Some works have explored tuning-free methods. ELITE Wei et al. (2023) and InstantBooth Shi et al. (2023) project reference images into word embeddings and inject reference image patch features into cross-attention layers to enhance local details. PhotoMaker Li et al. (2023b) focuses on the generation of human portraits by extracting a stacked ID embedding from multiple ID images. Despite impressive results for single-object customization, their architecture design restricts their scalability to multiple subject settings. Models like Subject-Diffusion Ma et al. (2023) and FastComposer Xiao et al. (2023) are designed for multi-entity subject-driven generation. They enhance text conditioning within diffusion models by incorporating subject embeddings extracted from image encoders.Yet, a prevalent limitation of these approaches is their inclination to separate textual and visual guidance, thereby constraining the efficacy of joint modality integration. This issue is particularly pronounced when extending to scenarios with multiple entities.
@@
Multimodal Large Language Models (MLLMs) have significantly broadened the capabilities of language models to process various modalities, including images. These models inherently facilitate interleaved vision-language input, effectively handling multiple images. Notable works include M-VADER Weinbach et al. (2022), which aligns MLLM semantics with the diffusion image decoder via image-caption pair training. Additionally, models such as GILL Koh et al. (2023), Emu Sun et al. (2023), Emu2 Sun et al. (2023) and DreamLLM Dong et al. (2023) specialize in interleaved vision-language generation, aligning the output space of MLLMs with the diffusion image decoder by leveraging CLIP supervision or pre-training on multimodal datasets. However, these methods primarily align at a semantic level and may struggle with detailed, subject-driven image generation. BLIP-Diffusion Li et al. (2023a) synthesizes images by composing subjects with random backgrounds, endowing it with zero-shot, subject-driven text-to-image generation capabilities. However, its specific input template and training process limit scalability for multiple entities. KOSMOS-G Pan et al. (2023), a model closely related to our work, leverages a MLLM to encode interleaved text-visual inputs. It utilizes the U-Net of Stable Diffusion (SD) v1.5 as its image decoder, which remains frozen throughout the training process. The key component of KOSMOS-G is an AlignerNet, trained solely on textual data, to align the output embedding space of the frozen SDv1.5 U-Net with the MLLM. In contrast, our approach centers on training the U-Net model end-to-end specifically for multimodal diffusion, significantly enhancing both fidelity and relevance of generated images in multimodal contexts. Differing from embedding alignment-based methodologies, our two-stage training strategy markedly improves the model’s capability in following multimodal instructions, which results in superior performance in both text-driven and subject-driven generation tasks, particularly in complex multi-entity scenarios.
@@ 4.1 Implementation Details
UNIMO-G is composed of a 7B-parameter MLLM encoder, following the MiniGPT-4 architecture Zhu et al. (2023), and a 4B-parameter denoising U-Net, totaling approximately 11B parameters. The MLLM is pretrained on a large-scale Chinese multimodal corpus, comprising text, image-caption pairs, and interleaved image-text data. The U-Net architecture includes 5 downsampling and 5 upsampling blocks with channel sizes [640, 1280, 2560, 2560], and a cross-attention mechanism with 4096 dimensions and 16 heads. The image auto-encoder, based on the LDM framework, has been optimized for our specific image corpus. The detail training process and data construction are further introduced in Appendix A.
@@
Figure 4: Comparison of UNIMO-G and SD-XL by human evaluation. The mean and standard deviation are shown in the figure.
4.2 Evaluation Results
Text-to-Image Generation
For text-to-image generation, we used 30,000 captions randomly sampled from the MS-COCO Lin et al. (2014) validation set, translating them into Chinese to align with UNIMO-G’s input requirements. Images were generated at 512x512 pixels and resized to 256x256 for evaluation using the FID-30k metric, a standard in the field. Our model employs a classifier-free guidance scale of 5.0 and 50 DDIM inference steps for diffusion sampling. As shown in Table 1, UNIMO-G significantly surpasses other Vision-Language to Image (VL2I) models in performance.
To further validate our model, we conducted a human evaluation comparing UNIMO-G with SD-XL Podell et al. (2023), a leading open-source model. We established a comprehensive bilingual benchmark, encompassing 312 prompts (162 from DrawBench and 150 user queries randomly sampled from online ***https://yige.baidu.com/ platforms). The DrawBench prompts were filtered to exclude language-specific ones. All prompts are manually translated and carefully proofread to achieve the final parallel Chinese and English set. Three independent evaluators rated the images from UNIMO-G and SD-XL by selecting the model they prefer, focusing on aspects of image aesthetics, image-text relevance, and overall quality, respectively. The results demonstrate UNIMO-G’s substantial superiority in all aspects.
@@ 4.2 Evaluation Results
Table 2: Comparisons of single-entity subject-driven image generation on DreamBench. Avg denotes the average score of DINO, CLIP-I and CLIP-T.
Single-Entity Subject-Driven Generation
Multi-Entity Subject-Driven Generation
UNIMO-G exhibits exceptional performance in zero-shot multi-entity subject-driven generation. To evaluate this capability, we established MultiBench, a novel benchmark specifically designed for multi-entity subject-driven generation assessment. MultiBench includes four object categories: living objects (humans and animals), food, wearable items, and toys, each containing 10 different objects. We developed five prompt templates for composing scenarios with 2 and 3 objects, resulting in a total of 2,300 distinct prompts. Details are introduced in the Appendix B. For each prompt, four images were generated, culminating in 9,200 images for an exhaustive evaluation. We conducted image similarity analysis using DINO and CLIP-I metrics, alongside text relevance assessments using CLIP-T. Image similarity was determined by averaging the similarities between the generated image and each of the two or three subjects. The results, as shown in Table 3, indicate that UNIMO-G outperforms BLIP-Diffusion and KOSMOS-G in terms of both image similarity and textual relevance. Some comparison examples are shown in Figure 5. This demonstrates UNIMO-G’s superior capability to accurately capture subject information from input images and effectively follow multi-modal instructions. More examples are shown in Figures 8 and 9.
@@ 4.3 Analysis
UNIMO-G, prior to multi-modal instruction tuning (denoted as “w/o Tuning”), also demonstrates the capability to generate images from multi-modal prompts based on the MLLM. Nonetheless, it initially falls short in accurately reproducing input images. To evaluate the effectiveness of multi-modal instruction tuning, we compared the performance of UNIMO-G with and without this tuning in single-entity and multi-entity subject-driven generation tasks. The comparison results in Table 2 and Table 3 reveal that multi-modal instruction tuning substantially enhances image similarity metrics (DINO and CLIP-I) in both single and multi-entity scenarios. This improvement indicates that after tuning, UNIMO-G more effectively leverages visual features from the inputs, thereby accurately replicating the content defined in the image conditions. Furthermore, UNIMO-G with instruction tuning also shows obvious advancements in textual relevance, as evaluated by the CLIP-T metric. This indicates that the tuning process not only bolsters image fidelity but also amplifies the model’s ability to follow multimodal instructions.

### v3

@@ Abstract
**footnotetext: These authors contributed equally to this work.footnotetext: https://unimo-ptm.github.io/
@@ 1 Introduction
Recent advancements in text-to-image (T2I) diffusion models have yielded impressive results in the generation of high-fidelity images from textual descriptions. Various methods, including DALL-Es Ramesh et al. (2022); Betker et al. (2023), Imagen Saharia et al. (2022), Stable Diffusion Rombach et al. (2022), and MM-DiT Esser et al. (2024), have been successful in producing photorealistic and contextually relevant images based on textual prompts. Nevertheless, a fundamental challenge persists due to the inherent brevity of textual descriptions, particularly when intricate details, specific entities, or nuanced scenes are involved. Thus, faithfully generating images from general vision-language (VL) inputs is essential to improve the controllability of image generation.
@@
UNIMO-G exhibits a comprehensive capability for controllable image generation, excelling not only in text-to-image synthesis but also in zero-shot subject-driven generation. It adeptly produces high-fidelity images from multimodal prompts, even those containing multiple image entities. To assess its performance, we conducted evaluations in both text-to-image and subject-driven generation contexts using the MS-COCO Lin et al. (2014) and DreamBench Ruiz et al. (2023) datasets, respectively. The results consistently highlight UNIMO-G’s superior performance in these scenarios. Additionally, recognizing DreamBench’s focus on single-subject generation, we introduce MultiBench, a new benchmark featuring images with multiple entities. The evaluation on MultiBench confirms UNIMO-G’s effectiveness in zero-shot multi-entity subject-driven generation.
@@ 2 Method
The architecture of UNIMO-G, as depicted in Figure 2, primarily comprises two key components: a Multimodal Large Language Model (MLLM) responsible for encoding multimodal prompts and a conditional denoising diffusion network for image generation based on the encoded representations. In our study, we employed a pre-trained Chinese MLLM, structurally similar to MiniGPT-4 Zhu et al. (2023). It consists of a vision encoder with a pretrained ViT and Q-Former, a single linear projection layer, and a Transformer-based LLM, underwent pre-training on a vast dataset comprising billions of image-text pairs. Details can refer to Appendix A. This extensive training process equips the model with a robust capability to process and interpret complex multimodal data.
Figure 2: UNIMO-G consists of an MLLM for multimodal perception, and a conditional denoising UNet for image generation. It accepts multimodal prompts with interleaved images and texts, and generates images consistent with the image entities. Orange denotes the trainable modules; Blue denotes the frozen ones.
The training of UNIMO-G is conducted in a two-stage process:
@@ 2.1 Text-to-Image Pre-training
$Attn(z_{t},c)=softmax(\frac{Q(z_{t})\cdot K(c)^{T}}{\sqrt{d}})\cdot V(c)$
@@
Training a text-to-image diffusion model from scratch presents significant challenges in terms of complexity and resource expenditure. To address these, we introduce an effective training schedule to enhance the efficiency and performance of model training. This schedule encompasses three phases: (1) initial training on a small image corpus to establish foundational visual distribution knowledge; (2) subsequent expansion to large-scale text-image pair training, focusing on text-visual correspondence; (3) culminating in a final phase of training with a small refined corpus, characterized by high visual aesthetics and precise text-image alignment. In our experiments, the training of the UNet model is initially conducted using the CC3M dataset Sharma et al. (2018). This dataset is chosen for its diverse range of visual concepts coupled with straightforward image descriptions, making it an effective tool for initiating training from scratch. Subsequently, the model undergoes further training with an extensive collection of 300M text-image pairs, aimed at broadening its conceptual understanding and improving its alignment to textual descriptions. The final phase of training involves fine-tuning the model using a meticulously curated corpus, consisting of tens of thousands of high-quality image-text pairs, carefully selected for their superior quality. Based on the above strategies and our architecture designs, we obtain a powerful Chinese text-to-image generation model, surpassing open-source models like Stable Diffusion and its advanced version SDXL Podell et al. (2023). We provide detailed evaluations of our results in Section 4.2 and implementation details in Appendix B.
@@ 2.2 Multimodal Instruction Tuning
To enhance the representativeness of text prompts, we introduce a format for multimodal prompts that are composed of interleaved images and texts. Specifically, entities mentioned in text captions can be substituted with their corresponding images, like “<img>image of Elon Musk</img> holding his <img>image of iPhone</img>, takes a selfie on the street”, as shown in Figure 2. To create pairs of multimodal prompts and images, we have designed a data processing pipeline as illustrated in Figure 3. The pipeline first generates captions and extracts entities from the caption by prompting the MLLM. Subsequently, it acquires the corresponding image segment for each entity using a combination of language grounding by Grounding DINO Liu et al. (2023b) and image segmentation by SAM Kirillov et al. (2023). Further introduction on the data construction process is provided in Section B. With a collection of pairs of multimodal prompts and images, UNIMO-G is trained to generate images in accordance with these multimodal instructions.
@@
$CA(z_{t},v)=Softmax(\frac{Q(z_{t})\cdot K(v)^{T}}{{\sqrt{d}}})$
@@ 3 Related Work
The incorporation of diffusion models into text-to-image synthesis represents a notable advancement in computational creativity Ho et al. (2020); Song et al. (2020); Li et al. (2022b). Models like GLIDE Nichol et al. (2021) and DALL-E 2 Ramesh et al. (2022), which utilize CLIP image embeddings, have substantially progressed in producing images that are both diverse and semantically coherent with textual inputs. Imagen Saharia et al. (2022) underscores the importance of language comprehension, proposing the integration of a large T5 language model to enhance semantic representation. The Latent Diffusion Model Rombach et al. (2022) addresses computational constraints by generating images from text-conditioned, low-dimensional latent spaces. Our proposed framework builds upon the principles of the Latent Diffusion Model, leveraging its computational efficiency and scalability.
@@
Following the success of generating high quality images from text descriptions, recent studies have explored subject-driven generation techniques. Models like DreamBooth Ruiz et al. (2023), textual-inversion Gal et al. (2022), and custom-diffusion Kumari et al. (2023) use optimization-based methods to embed subjects into diffusion models. This is achieved by either fine-tuning the model weights or inverting the subject image into a text token that encodes the subject identity. Some works have explored tuning-free methods. ELITE Wei et al. (2023) and InstantBooth Shi et al. (2023) project reference images into word embeddings and inject reference image patch features into cross-attention layers to enhance local details. PhotoMaker Li et al. (2023b) focuses on the generation of human portraits by extracting a stacked ID embedding from multiple ID images. Despite impressive results for single-object customization, their architecture design restricts their scalability to multiple subject settings. Models like Subject-Diffusion Ma et al. (2023) and FastComposer Xiao et al. (2023) are designed for multi-entity subject-driven generation. They enhance text conditioning within diffusion models by incorporating subject embeddings extracted from separate image encoders.Yet, a prevalent limitation of these approaches is their inclination to separate textual and visual guidance, thereby constraining the efficacy of joint modality integration.
@@
Multimodal Large Language Models (MLLMs) have significantly broadened the capabilities of language models to process various modalities Liu et al. (2023a); Li et al. (2021, 2022a); Wang et al. (2023); Driess et al. (2023). These models inherently facilitate interleaved vision-language input, effectively handling multiple images. Models such as GILL Koh et al. (2023), Emu Sun et al. (2023), and DreamLLM Dong et al. (2023) specialize in interleaved vision-language generation by aligning the output space of MLLMs with the diffusion image decoder. However, these methods primarily align at a semantic level and may struggle with detailed, subject-driven image generation. BLIP-Diffusion Li et al. (2023a) synthesizes images by composing subjects with random backgrounds, endowing it with zero-shot, subject-driven text-to-image generation capabilities. However, its specific input template and training process limit scalability for multiple entities. KOSMOS-G Pan et al. (2023), a model closely related to our work, leverages a MLLM to encode interleaved text-visual inputs, and the U-Net of Stable Diffusion (SD) v1.5 as the image decoder. The key component of KOSMOS-G is an AlignerNet, which is trained solely on textual data to align the output embedding space of the frozen SDv1.5 U-Net with the MLLM. In contrast, our approach centers on training the U-Net model end-to-end specifically for multimodal diffusion, significantly enhancing both the faithfulness and relevance of generated images in multimodal contexts. Differing from alignment-based approaches, our two-stage training strategy markedly improves the model’s proficiency in following multimodal instructions, particularly in complex multi-entity scenarios.
@@ 4.1 Implementation Details
UNIMO-G is composed of a 7.8B-parameter MLLM encoder, following the MiniGPT-4 architecture Zhu et al. (2023), and a 4B-parameter denoising U-Net, totaling approximately 11.8B parameters. The MLLM is pretrained on a large-scale Chinese multimodal corpus, comprising text, image-caption pairs, and interleaved image-text data. The U-Net architecture includes 5 downsampling and 5 upsampling blocks with channel sizes [640, 1280, 1280, 2560, 2560], and a cross-attention mechanism with 4096 dimensions and 16 heads. The image auto-encoder, based on the LDM framework, has been optimized for our specific image corpus. The detail training process and data construction are further introduced in Appendix B.
4.2 Evaluation Results
Text-to-Image Generation
For text-to-image generation, we used 30,000 captions randomly sampled from the MS-COCO Lin et al. (2014) validation set, translating them into Chinese to align with UNIMO-G’s input requirements. Images were generated at 512x512 pixels and resized to 256x256 for evaluation using the FID-30k metric, a standard in the field. Our model employs a classifier-free guidance scale of 5.0 and 50 DDIM inference steps for diffusion sampling. As shown in Table 1, UNIMO-G greatly surpasses other Vision-Language to Image (VL2I) models in performance.
To further validate our model, we conducted a human evaluation comparing UNIMO-G with SDXL Podell et al. (2023), a leading open-source model. We established a comprehensive bilingual benchmark, encompassing 312 prompts (162 from DrawBench and 150 user queries randomly sampled from the online platformhttps://yige.baidu.com/). The DrawBench prompts were filtered to exclude language-specific ones. All prompts are manually translated and carefully proofread to achieve the final parallel Chinese and English set. Three independent evaluators rated the images from UNIMO-G and SDXL by selecting the model they prefer, focusing on aspects of image aesthetics, image-text relevance, and overall quality, respectively. The results in Figure 4 demonstrate UNIMO-G’s substantial superiority in all aspects. Some examples are shown in Figure 11.
@@ 4.2 Evaluation Results
Single-Entity Subject-Driven Generation
Figure 4: Comparison of UNIMO-G and SDXL by human evaluation. The mean and standard deviation are shown in the figure.
Multi-Entity Subject-Driven Generation
UNIMO-G exhibits exceptional performance in zero-shot multi-entity subject-driven generation. To evaluate this capability, we established MultiBench, a novel benchmark specifically designed for multi-entity subject-driven generation assessment. MultiBench includes four object categories: living objects (humans and animals), food, wearable items, and toys, each containing 10 different objects. We developed five prompt templates for composing scenarios with 2 and 3 objects, resulting in a total of 2,300 distinct prompts. Details are introduced in the Appendix D. For each prompt, four images were generated, culminating in 9,200 images for an exhaustive evaluation. We conducted image similarity analysis using DINO and CLIP-I metrics, alongside text relevance assessments using CLIP-T. Image similarity was determined by averaging the similarities between the generated image and each of the two or three subjects. The results, as shown in Table 3, indicate that UNIMO-G outperforms BLIP-Diffusion and KOSMOS-G in terms of both image similarity and textual relevance. Some comparison examples are shown in Figure 6. This demonstrates UNIMO-G’s superior capability to accurately capture subject information from input images and effectively follow multi-modal instructions. More examples are shown in Figures 9 and 10.
@@
Table 2: Comparisons of single-entity subject-driven image generation on DreamBench. Avg denotes the average scores of DINO, CLIP-I and CLIP-T.
@@ 4.3 Analysis
UNIMO-G, prior to multi-modal instruction tuning (denoted as “w/o Tuning”), also demonstrates the capability to generate images from multi-modal prompts based on the MLLM. Nonetheless, it initially falls short in accurately reproducing input images. To evaluate the effectiveness of multi-modal instruction tuning, we compared the performance of UNIMO-G with and without this tuning in single-entity and multi-entity subject-driven generation tasks. The comparison results in Table 2 and Table 3 reveal that multi-modal instruction tuning substantially enhances image similarity metrics (DINO and CLIP-I) in both single and multi-entity scenarios. This improvement indicates that after tuning, UNIMO-G more effectively leverages visual features from the inputs, thereby accurately replicating the content defined in the image conditions. Furthermore, UNIMO-G with instruction tuning also shows obvious advancements in textual relevance, as evaluated by the CLIP-T metric. This indicates that the tuning process not only bolsters visual faithfulness but also amplifies the model’s ability to follow multimodal instructions.